The Coronavirus Disease is an infectious disease caused by a new strain of coronavirus. The World Health Organization claims that the incubation period of the virus in the infected person has a mean of 6.1 days. The doctors in the Philippines conducted a research and they found out that incubation period of the virus in human body is 7.03 days with a standard deviation of 4.32. The samples were 46 COVID patients. Is there enough evidence to conclude that the incubation period of the virus is 6.1 days as stated, at a=0.01?



1
Expert's answer
2021-06-17T11:29:43-0400

"H0:\mu=6.1"

"Ha:\mu\ne6.1"

Population variance is unknown, thus, t-distribution is used.

"t=\frac{\bar X-\mu}{\frac{s}{\sqrt{n}}}"

"t=\frac{7.03-6.1}{\frac{4.32}{\sqrt{46}}}=1.46"

"Cv=t_{0.005,45}=2.69"

Since 1.46<2.69, we fail to reject the null hypothesis and conclude that there is sufficient evidence to support the claim that the incubation period for the virus is 6.1 days.